OpenStudy (anonymous):

which following is[5-6i] a.iroot 6i b. root 11 c.root 61 d.root 11i

OpenStudy (anonymous):

option c

OpenStudy (anonymous):

\[\left| 5-6i \right|=\sqrt{5^2+6^2}=\sqrt{61}\]

OpenStudy (anonymous):

this does mean absolute value yes?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then saubhik wrote it exactly right.

OpenStudy (anonymous):

it is just \[\sqrt{a^2+b^2}\] don't think about the i or the - sign. just square, add, take the square root

OpenStudy (anonymous):

ok thanks;)